OpenStudy (adalrymple):

Matthew filled his garden with 12 cubic feet of organic dirt. The garden is a rectangular prism with a volume V equals l times w times h where l is length, w is width, and h is depth. If the garden has a depth of 1.5 feet and is 2 feet wide, how long is the garden? Select one: A. 2 feet B. 3 feet C. 4 feet D. 5 feet

OpenStudy (isaidavila):

V=lwh so 12=l*2*1.5

OpenStudy (anonymous):

\[V=l \times w \times h\] therefore we can rearrange to get\[l=\frac{V}{w \times h}\] and then substitute known values to solve.

OpenStudy (adalrymple):

@isaidavila @twvogels I think I did something wrong. I got 8.5 but that's not one of the options

OpenStudy (isaidavila):

12=L*2*(1.5) 12=L*3 L=12/3

OpenStudy (adalrymple):

Ohhh ok thank you:) @isaidavila

OpenStudy (isaidavila):

No problem